Circular No. 7342

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gr IN MCG +5-29-24 G. M. Hurst, Basingstoke, England, reports the discovery by Mark

Armstrong, Rolvenden, of an apparent supernova on unfiltered CCD images taken on 1999 Dec. 13 and 15 with a 0.30-m telescope in the course of the U.K. Nova/Supernova Patrol. SN 1999gr is located at α = 12^{h}09^{m}11^{s}.31, δ = +29^{o}16′34′′.7 (equinox 2000.0), which is 23′′ east and 8′′ north of the nucleus of MCG +5-29-24. Hurst reports the following approximate magnitude estimates from CCD images taken by Armstrong (Dec. 13 and 15) and by T. Boles, Wellingborough (Dec. 19): Dec. 13.185 UT, 18.5; 15.166, 18.6; 19.173, 18.6. A CCD image obtained by M. Schwartz, Cottage Grove, OR, obtained on 1999 Jan. 14 does not show the new object (limiting mag ∼ 18.5–19.0), and SN 1999gr is not visible on second Palomar Sky Survey images (limiting mag ∼ 20.5).

S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics; and P. Garnavich, University of Notre Dame, report that a spectrum of SN 1999gr, obtained by M. Calkins on Jan. 5.46 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova, several months after maximum light. Though of low signal-to-noise, the spectrum exhibits broad emission features of [Fe II] and [Fe III], observed at 470, 530, and 590 nm.

COMET C/1999 S4 (LINEAR) D. G. Schleicher, Lowell Observatory, writes: “Narrowband photom-

etry of comet C/1999 S4 was obtained with the Hall 1.07-m telescope at Lowell Observatory on 1999 Dec. 5, 28, and 30 UT (r = 3.57–3.28 AU). The only gas species definitively detected has been CN, with a near-constant Haser production rate of log Q(CN) = 24.2; C_{2} was marginally detected in late December. The upper limit on OH emission yields log Q(OH) < 27. The dust-production rates were log (Afρ) = 2.4 on Dec. 5 and 2.2 on Dec. 28 and 30 (A is the albedo of the grains, f is the filling factor of the grains, and ρ is the radius of the field-of-view; cf. A’Hearn et al. 1984, A.J. 89, 579). This apparent decrease in dust production implies either an earlier outburst or significant variability due to rotation. Should the comet be dynamically new, the decrease in dust production likely predicts behavior similar to that of comet C/1973 E1 (Kohoutek), due to the continued dispersion of dust released at much larger r.”

2000 January 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green